OpenStudy (vinicius1):

Limit sin(1/x) as x->0 diverges, right? so is it continuous where x->0?

OpenStudy (amistre64):

continuous implies something about the value at x=0

OpenStudy (amistre64):

its continuous as it approaches zero, yeah .... it just alternates between -1 and 1 and never settles between the two

OpenStudy (vinicius1):

Thanks. I think there's a mistake in the question then. It says that it is continuous where x=0, but when x=0 then its undefined, right?

OpenStudy (amistre64):

1/x is undefined at x=0 and as such, there is no value to attribute to the sine of an undefinable angle

OpenStudy (amistre64):

the limit of a function doesnt depend on the value of a function at x, it depends on if you can get the same value by aproaching it from all directions

OpenStudy (amistre64):

is the function continuous as it approaches 1/0? sure, for sin is defined for all values that are definable

OpenStudy (vinicius1):

the question is actually this one: Prove that the function f(x)=x(sin(1/x), when x is not 0, f(x)=0 when x=0, is continuous at x=0 but isnt differentiable at the same point

OpenStudy (vinicius1):

sorry for poor translation

OpenStudy (amistre64):

so we seem to have a squeeze thrm here

OpenStudy (amistre64):

the amplitude of the function goes to zero, so the sin function will fluctuate between the lines y=x and y=-x which of course funnels in at x=0

OpenStudy (amistre64):

differentiable, id assume we just take a derivative and see what it results

OpenStudy (amistre64):

not sure what methods are availble to you

OpenStudy (vinicius1):

In this case I was told to use the definition of the derivative. It wasn't differentiable

OpenStudy (amistre64):

sin(u) = u - u^3/3 + u^5/5 - u^7/7 + ... sin(1/x) = 1/x - 1/3x^3 + 1/5x^5 - 1/7x^7 + ... xsin(1/x) = 1 - 1/3x^2 + 1/5x^4 - 1/7x^6 + ... that would be the taylor polynomial for it if that helps lol

OpenStudy (vinicius1):

Yes, it does. But the tutor told me to use the definition f'(x)=lim (f(x+h) - f(x))/h as h->0 and doing that i end up with lim sin(1/x) as x->0

OpenStudy (amistre64):

i forgot those were factorials ... not digits

OpenStudy (vinicius1):

Dont worry

OpenStudy (amistre64):

x sin(1/x) x' sin(1/x) + xcos(1/x) (-1/x^2) sin(1/x) - [cos(1/x)] / x the derivative would get us that

OpenStudy (amistre64):

which is not definable at x=0

OpenStudy (vinicius1):

so it is not continuous at x=0?

OpenStudy (zarkon):

are you asking if \[\sin\left(\frac{1}{x}\right)\] is continuous at 0?

OpenStudy (amistre64):

you keep asking for continuous, while speaking of limits

OpenStudy (amistre64):

if the limit of f(x), as x approaches c+ and c-, is equal to f(c) then it is continuous at x=c if memory serves for a definition of continuity :)

OpenStudy (vinicius1):

The question asks to prove that the function f(x)= x(sin(1/x)) for x is not 0, and f(x)=0 for x=0, is continuous but not differentiable. so its assuming its continuous

OpenStudy (vinicius1):

continuous at x=0 but not diff. at the same point

OpenStudy (vinicius1):

It's the fourth question. translation "prove that the function f(x) is continuous at x=0 but not differentiable at the same point"

OpenStudy (amistre64):

ah, you have a piecewise function

OpenStudy (amistre64):

by the squeeze thrm, the limit of f(x) as x approaches 0 from the left and right is ... 0 and by definition of f(x), f(0)=0 since the limit and the function are the same at x=0, then it is continuous at x=0 by the crude recollection of a definition of continuity

OpenStudy (vinicius1):

hmm that makes sense now. but it's also not differentiable, right?

OpenStudy (vinicius1):

at x=0

OpenStudy (amistre64):

derivatives deal with slopes, can we conclude what the slope of the line is at x=0? that is where the limit definition would prolly be applied

OpenStudy (amistre64):

its like trying to find the derivative of |x| at x=0

OpenStudy (amistre64):

at any rate ... sin(1/x) - [cos(1/x)]/x is our derivative and successive derivatives are just as absurd. the taylor polynomial helps to see this better

OpenStudy (amistre64):

helps ME to see this better that is :)

OpenStudy (zarkon):

just apply the definition of the derivative you will end up with computing the limit\[\lim_{h\to 0}\sin\left(\frac{1}{h}\right)\] which is easy to do

OpenStudy (vinicius1):

I see. Look, this is the answer they told me is the right one: hmm. lim (h -> 0) f(h) / h = lim (h -> 0) h*sen(1/h)/h so lim (h->0) sen(1/h) 1/h -> infinity when h->0 lim (x-> infinity) sen(x) does not exist, so its not differentiable at x=0

OpenStudy (amistre64):

looks like zarkons reply :)

OpenStudy (vinicius1):

hmm makes sense now. so lim sin(infinity) does not exist?

OpenStudy (zarkon):

the sine function if two pi periodic. so you just keep getting the same stuff over and over as x goes to infinity. it never converges

OpenStudy (vinicius1):

It diverges, right? I just don't get how it is continuous if it never converges. I'm kind of new to this.

OpenStudy (amistre64):

can you determine the slope of |x| at x=0? continuity and limit are 2 different things.

OpenStudy (amistre64):

as well as slope

OpenStudy (vinicius1):

hmm. I understand. Thanks to you all.

OpenStudy (amistre64):

yw